OpenStudy (anonymous):

is 6/24 a ratinol or irratinol

OpenStudy (anonymous):

... 6/24 = 1/4 = fraction = rational.

OpenStudy (anonymous):

is 10/33 a ratinol or irratinol

OpenStudy (anonymous):

. fraction. what do you think?

OpenStudy (anonymous):

umm irratinol lol jk

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

is \[\div\]18 a ratinol or iratinol #?

OpenStudy (anonymous):

......fractions are things with a decimal sign. i refer to you: http://www.visualfractions.com/

OpenStudy (anonymous):

im not going there and this has no decimal sighn

OpenStudy (anonymous):

..............÷ is a decimal sign as well. it is just another way of writing it. sort of like \[\times\] and *